# Route to Chaos and Unified Dynamical Framework of Multi-Species Ecosystems

### v1

@@ Abstract
We investigate species-rich mathematical models of ecosystems. Much of the existing literature focuses on the properties of equilibrium fixed points, in particular their stability and feasibility. Here we emphasize the emergence of limit cycles following Hopf bifurcations tuned by the variability of interspecies interaction. As the variability increases, and owing to the large dimensionality of the system, limit cycles typically acquire a growing spectrum of frequencies. This often leads to the appearance of strange attractors, with a chaotic dynamics of species abundances characterized by a positive Lyapunov exponent. We find that limit cycles and strange attractors preserve biodiversity as they maintain dynamical stability without species extinction. We give numerical evidences that this route to chaos dominates in ecosystems with strong enough interactions and where predator-prey behavior dominates over competition and mutualism. Based on arguments from random matrix theory, we further conjecture that this scenario is generic in ecosystems with large number of species, and identify the key parameters driving it. Overall, our work proposes a unifying framework, where a wide range of population dynamics emerge from a single model.
@@
Introduction. One of the main challenges in theoretical ecology is to connect predictions from mathematical models of population dynamics to empirical observations of species coexistence in natural or laboratory-controlled ecosystems May07 . It is established that individual populations fluctuate in time, often with large qualitative and quantitative differences between species Ran98 ; Lun00 ; Inc03 ; Van06 . In some instances, population abundances exhibit synchronized, periodic oscillations, while in others population dynamics appears chaotic Has93 ; Ell95 ; Hun98 ; Hui99 ; Ben08 ; Rog21 ; Pre10 ; Pea20 . It is highly desirable to find out whether the wide range of observed dynamics can be captured by varying few parameters of a unifying mathematical model, and if yes, what are the key characteristics this model should retain. Investigating mathematical models can furthermore shed light on fundamental qualitative questions such as whether populations fluctuations are endogenous or exogenous, i.e., if they are generated by intrinsic interactions or by external sources Ell95 ; Sch03 .
@@
In this letter, we numerically investigate large Lotka-Volterra models with random interactions, which are standard multi-species models of population dynamics. We emphasize the richness of their dynamics as a function of two key parameters which are (i) the variability $\sigma$ in interspecies interactions, and (ii) the off-diagonal covariance parameter $\gamma$ of the interaction matrix. Our main finding is that, for sufficiently large ratio of predator-prey pairs of species, the stable fixed-points prevailing at weak interaction variability generically lose their stability through Hopf bifurcations. Limit cycles emerge, where surviving species have periodically oscillating abundances. At still stronger interactions, strange attractors appear, possibly from cascades of bifurcations, which lead to a chaotic dynamics of population abundances characterized by a positive largest Lyapunov exponent. This route to chaos illustrates how stationarity, oscillating periodicity and chaos in the dynamics of species abundances exist in rather general models of population dynamics, depending on $\sigma$ and $\gamma$. One important result is that all observed population dynamics in multi-species ecosystems can be reproduced by a unified mathematical model.
@@
Much of theoretical ecology is based on the surmise that the observed states and dynamics of ecosystems can be described by the time-asymptotic behavior of mathematical models. Accordingly, there has been a significant focus on the equilibrium fixed points of large systems of competing species, their stability and feasibility May72 ; Rob74 ; Rie89 ; All12 ; Tan14 ; Fyo16 ; Cle23 ; Ros23 One key parameter is the variability $\sigma$ of the interspecies interaction. It has been found that a unique, asymptotically stable fixed point exists at weak enough $\sigma<\sigma_{c}$ May72 ; Rie89 ; Cle23 ; Ros23 , and that $\sigma_{c}$ increases in ecosystems with dominating predator-prey interactions All12 ; Tan14 . For $\sigma>\sigma_{c}$, one enters a phase with multiple unstable equilibria Ros23 . Stabilization may still occur via species extinction, which drives the ecosystem to a novel, stable fixed point with reduced biodiversity. Beyond fixed points, the emergence of limit cycles through Hopf bifurcation has been emphasized, albeit for small systems with only few species Gil75 ; Cos79 ; Gar89 ; Fus00 ; Mcg08 . Further regimes with aperiodic persistent dynamics have recently been highlighted at stronger interactions and/or larger number of species Hu22 , in ecosystems with sparse interactions Roy20 ; Mar24 or communities with migrations Dal20 . In contrast to this mainstream philosophy focusing on the time-asymptotic behavior of mathematical models, several works have advocated that sufficiently large ecosystems have long relaxation times so that, by nature, they are in a transient state Has94 ; Schr03 ; Mor16 ; Has18 . Such large relaxation times may follow from the slowing down of the dynamics in ecosystems close to criticality Sol02 ; Mor11 . Below we show that all these different behaviors naturally emerge in different regimes of a single mathematical model, as a function of only two parameters.
@@
Model and method. Population dynamics in multi-species ecosystems is commonly studied in the framework of the generalized Lotka-Volterra model Goe71
@@
Eq. (1) determines the time-evolution of the normalized abundance $N_{i}(t)\geq 0$ of species $i=1,2,\ldots S$, as a function of its intrinsic growth rate $\kappa_{i}$ and its interaction with other species. Interactions have a finite average $\mu$ and a variability $\sigma$. Fluctuations in interaction strengths between pairs of species are encoded in the components ${\mathbb{A}}_{ij}$ of the interaction matrix $\mathbb{A}$. Being interested in large, heterogeneous ecosystems with no particular structure, we follow a random matrix theory (RMT) approach May72 ; All15 ; Akj24 and take ${\mathbb{A}}_{ij}$ to be normally distributed with vanishing average and covariances given by
@@
The off-diagonal covariance parameter $\gamma\in[-1,1]$ tunes from competitive or mutualistic pairs of species ($\gamma=1$; where the spectrum of $\mathbb{A}$ is real) to only predator-prey pairs of species ($\gamma=-1$; ; where the spectrum is purely imaginary) gamma . For $\gamma=0$, $\mathbb{A}$ belongs to the Ginibre ensemble or random matrices Gin65 .
The generalized Lotka-Volterra model of Eq. (1) assumes that $N_{i}(t)$ is real and varies continuously. This is a legitimate assumption only as long as $N_{i}(t)$ is sufficiently large. Large fluctuations have been observed in numerical simulations of Eq. (1), where some species resurrect from minuscule abundances, effectively corresponding to extinction. As a matter of fact, the dynamics of Eq. (1) leads to species extinction only for asymptotically long times, and the standard procedure to solve this atto-fox problem Mol91 is to introduce a small, but finite extinction threshold $N_{c}$. When $N_{i}(t_{c})<N_{c}$, extinction occurs and $N_{i}(t>t_{c})\equiv 0$. We will set $N_{c}=10^{-20}$ but have checked that other choices lead to the same conclusions as presented below. Our focus is on the parameters $\sigma$ and $\gamma$ defining the variability of the interspecies interactions and accordingly we fix the growth rates, the initial number of species and the average interaction at values $\kappa_{i}\equiv 1$, $S=157$ and $\mu=5$. We have checked that varying $\mu$ does not change our conclusions as long as $\mu>0$.
@@
The RMT approach to ecosystem dynamics dates back at least to May’s seminal work May72 . Fixed-points $\vec{N^{*}}$ of Eq. (1) are defined by $\dot{N}_{i}^{*}=0$ and the dynamics in their vicinity is governed by the stability matrix $\mathbb{M}$
@@
$\mathbb{M}_{ij}$ $=$ $-N_{i}^{*}(\delta_{ij}+\mu/S+\sigma{\mathbb{A}}_{ij}/\sqrt{S})\,.$ (3b)
The fixed-point is stable as long as the spectrum of $\mathbb{M}$ lies entirely in the left half of the complex plane. The average density of eigenvalues of $\mathbb{A}$ is distributed within a zero-centred ellipse in the complex plane with semi-axes $\sigma(1+\gamma)$ [$\sigma(1-\gamma)$] in the real [imaginary] direction Som88 . The ellipse is shifted to the left by the identity matrix [the Kronecker symbol in Eq. (3b)], while the constant $\mu$-term in Eq. (3b) adds an outlier eigenvalue $-\mu$ Cle23 . Assuming an homogeneous distribution of populations, $N_{i}^{*}\simeq N_{0}={\cal O}(1)$, the fixed-point is parametrically stable for $\sigma<(1+\gamma)^{-1}$. Beyond that border, fixed-point stability may be recovered via species extinctions, $N_{i}^{*}\rightarrow 0$, except for a number $N_{s}<S$ of surviving species, because then, the stability matrix $\mathbb{A}^{(r)}$ is reduced by removing rows and lines from $\mathbb{A}$, corresponding to the extinct species. Then, the spectrum of $\mathbb{A}^{(r)}$ has an elliptic support with semi-axes $\sigma(1\pm\gamma)\sqrt{N_{s}/S}$. Stability then bounds the number of surviving species $N_{s}/S\leq[\sigma(1+\gamma)]^{-2}$. Note that this argument says nothing about the bifurcation through which fixed-points lose stability.
@@
Results. For $\gamma\neq 1$ the eigenvalues of the real, asymmetric matrix $\mathbb{M}$ are either real, or come in complex-conjugated pairs. This opens up the possibility that the fixed-point loses its stability through a Hopf bifurcation, after which the ecosystem dynamics is attracted to a stable limit cycle. Such a Hopf bifurcation is illustrated in Fig. 1, where the top panel shows periodic oscillations in abundances of the surviving species which are related in the bottom panel to the crossing of the imaginary axis by a complex-conjugated pair of eigenvalues. For large random matrices with $\gamma=0$ it has been shown that only ${\cal O}(\sqrt{S})$ of the $S$ eigenvalues are real Ede94 ; Ede97 ; Kan05 , which suggests that such Hopf bifurcations are the rule rather than the exception in ecosystems with high biodiversity.
Directly following the bifurcation, cycles exhibit sinusoidal oscillations with a single frequency, determined by the imaginary part of the pair of involved eigenvalues. More harmonics emerge as $\sigma$ increases further, until the cycle disappears. Cycle disapearance can happen because the cycle loses either its stability – for instance through an inverse Hopf bifurcation – or its feasibility – for instance because one or several species reach the extinction threshold $N_{c}$ somewhere along the cycle. The dynamics of a single-species as $\sigma$ passes through and keeps increasing beyond a Hopf bifurcation is shown in Fig. 2. It illustrates a third mechanism by which a periodic cycle turns into a strange attractor. To quantify this transition to chaos, we numerically calculated the largest Lyapunov exponent $\lambda$ Ben80 . As expected, $\lambda<0$ in the fixed-point regime, $\sigma\lesssim 2.59$. Following the Hopf bifurcation, $\lambda=0$ as long as the limit cycle remains stable [panels (b) and (c)], which corresponds to the dynamics in the direction tangential to the cycle. Upon further increase of $\sigma$, $\lambda>0$ as one enters the chaotic regime, with a population dynamics governed by a strange attractor. The black and red trajectories in panels (d) and (e) illustrate the associated sensitivity to initial conditions, where population trajectories repeat similar-looking patterns, albeit at irregular time intervals and following sequences depending strongly on initial conditions. We stress that the abrupt fluctuations exhibited by $\lambda$ in both the fixed-point and the strange attractor regimes reflect fast dynamical changes with $\sigma$. Numerical error bars in panel (f) of Fig. 2 are smaller than symbol sizes, in particular, $|\lambda|\lesssim 10^{-5}$ in the limit cycle regime $2.59\lesssim\lambda\lesssim 2.82$ (grey area in Fig. 2).
@@
The occurence of Hopf bifurcations is of central importance for theoretical ecology. It extends the stability of population coexistence beyond the loss of fixed-point stability, without the need for species extinction. The resulting increasing biodiversity is illustrated in Fig. 3. It is obvious that the average number of coexisting species at fixed $\sigma$ is higher for ecosystems equilibrating to limit cycles ($\langle N_{s}\rangle=58.4$ and $63$) than to fixed-points ($\langle N_{s}\rangle=51.8$ and $53.7$). The prevalence of limit cycles further depends on both the variability $\sigma$ of interspecies interaction and on the off-diagonal covariance parameter $\gamma$. As a matter of fact, to have a limit cycle, one needs a large enough $\sigma$ to have a bifurcation in the first place. Second, the bifurcation must also involve a pair of complex-conjugated eigenvalues of the stability matrix. Both the value of $\sigma$ and the probability that a complex-conjugated pair of eigenvalues triggers the bifurcation depend on $\gamma$. Fig. 3 suggests that more negative values of $\gamma$, together with the associated larger value of $\sigma$ to have a bifurcation favor the occurence of limit cycles. This fact is confirmed in Fig. 4, which shows the probability $P(\gamma)$ that the ecosystem reaches a state of persistent dynamics. The numerical detection method employed for this statistical analysis does not differentiate between limit cycles and chaotic motion on a strange attractor, however sampling those data indicate that chaotic motion sets in only at larger interactions, and constitutes a significant fraction of the data only for $\sigma=4$. Work to better quantify this fraction is currently underway. Data for larger $\sigma$ are not shown as they exhibit mass extinctions, with only very few surviving species at best, except for the smallest values of $\gamma$.
Figure 4: Probability to find the system in a state of persistent dynamics – either a limit cycle or a state of chaotic dynamics on a strange attractor – as a function of $\gamma$, and for $\sigma=1.5$ [Panel(a)], 2 [(b)], 3 [(c)] and 4 [(d)]. Data have been calculated over ensembles of 100 different interaction matrices, each with 5 different initial conditions.
The observed increased probability to find limit cycles at negative values of $\gamma$ is related to the associated increased fraction of pairs of complex-conjugated eigenvalues in the spectrum of the stability matrix. Neglecting inhomogeneities in the fixed-point abundancies $N_{i}^{*}$, this spectrum is real for $\gamma=1$ and purely imaginary for $\gamma=-1$, and it is expectable that the fraction of real eigenvalues decreases as $\gamma$ decreases. To the best of our knowledge, the only mathematically rigorous result at intermediate values of $\gamma$ is that only a fraction ${\cal O}(\sqrt{S})$ of the eigenvalues are real for the Ginibre ensemble, i.e., at $\gamma=0$ Ede94 ; Ede97 ; Kan05 . While these observations suggest more frequent occurences of limit cycles as $\gamma$ decreases, what truly matters is whether the extreme eigenvalue with largest real part is real or complex. Calculating how likely that is as a function of $\gamma$ is a formidable task and we are unaware of any rigorous result in this direction. We therefore resort to numerical calculations (See Supplemental Material SM , Fig. S1). As expected, we find that the probability of a complex extreme eigenvalue increases with decreasing $\gamma$. To translate this result into a probability to have a limit cycle emerge through a Hopf bifurcation, we still need to take into account that there is no bifurcation at $\gamma=-1$, since there, increasing $\sigma$ only stretches the spectrum in the imaginary direction. We therefore expect that moderately negative values of $\gamma$ should favor the emergence of limit cycles. This qualitative argument is confirmed by the data shown in Fig. 4. A second numerical result shown in Fig. S1 SM is that this probability increases with increasing number $S$ of species.
@@
We translate these results into the language of theoretical ecology. Noting that $\gamma$ interpolates between ecosystems consisting purely of pairs of either mutualistic or competitive species for $\gamma=1$, to ecosystems with only predator-prey pairs of species for $\gamma=-1$ gamma , persistent oscillating behavior in population abundances are expected to be prevailing in large, sufficiently interacting ecosystems with a majority of predator-prey pairs.

### v2

# Route to Chaos in Multi-Species Ecosystems

@@ Abstract
We investigate species-rich mathematical models of ecosystems. While much of the existing literature focuses on the properties of equilibrium fixed points, persistent dynamics (e.g., limit cycles or chaos) have also been observed, both in natural or lab-controlled ecosystems and in mathematical models. Here we emphasize the emergence of limit cycles following Hopf bifurcations tuned by the variability of interspecies interaction. As this variability increases, and owing to the large dimensionality of the system, limit cycles typically acquire a growing spectrum of frequencies. This often leads to the appearance of strange attractors, with a chaotic dynamics of species abundances characterized by a positive Lyapunov exponent. We observe that limit cycles and strange attractors preserve biodiversity to some extent, as they maintain dynamical stability without species extinction. We give numerical evidences that this route to chaos dominates in ecosystems with strong enough interactions and where predator-prey behavior dominates over competition and mutualism. Based on arguments from random matrix theory, we further conjecture that this scenario is generic in ecosystems with large number of species, and identify the key parameters driving it. Overall, we show that the model we consider provides a unifying framework, where a wide range of population dynamics emerge from a simple few-parameter model.
The time evolution of a species population in a given ecosystem can follow very diverse paths. Of course, a population can remain relatively stable over time, but it can also oscillate between highs and lows, from one year to another. More surprisingly, recent studies have shown that this evolution can even be chaotic – meaning that no regular patterns emerge, and the population at any given time is highly sensitive to its initial state. We show that the three behaviors aforementioned (steady state, cycle, and chaos) can be modeled in a parcimonious version of the generalized Lotka-Volterra model. Furthermore, we identify the two model parameters that govern the transition between these three states. Namely, we observe that, as the interaction strengths between species increases, ecosystems tend to transition from a steady state, to a cyclic trajectory, and further to a chaotic behavior.
I Introduction
@@
One of the main challenges in theoretical ecology is to connect predictions from mathematical models of population dynamics to empirical observations of species coexistence in natural or laboratory-controlled ecosystems. May, Robert and McLean, Angela R (2007) In the latter it is established that individual populations fluctuate in time, often with large qualitative and quantitative differences between species. Ranta, Kaitala, and Lundberg (1998); Lundberg et al. (2000); Inchausti and Halley (2003); Vandermeer (2006) In some instances, population abundances exhibit synchronized, periodic oscillations, while in others population dynamics appears chaotic. Hastings et al. (1993); Ellner and Turchin (1995); Hunter and Price (1998); Huisman and Weissing (1999); Benincà et al. (2008); Rogers, Johnson, and Munch (2022); Prendergast et al. (2010); Pearce, Agarwala, and Fisher (2020) It is highly desirable to find out whether the wide range of observed dynamics can be captured in a single mathematical model, by varying few parameters, and if yes, what are the key characteristics this model should retain. Investigating mathematical models can furthermore shed light on fundamental qualitative questions such as whether populations fluctuations are endogenous or exogenous, i.e., if they are generated by intrinsic interactions or by external sources. Ellner and Turchin (1995); Scheffer et al. (2003)
@@
In this manuscript, we numerically investigate large Lotka-Volterra models with random interactions, which are standard multi-species models of population dynamics. We emphasize the richness of their dynamics as a function of two key parameters which are (i) the variability $\sigma$ of interspecies interactions, and (ii) the cross-diagonal covariance parameter $\gamma$ of the interaction matrix. Our main finding is that, for a predominance of predator-prey pairs of species, the stable fixed-points prevailing at weak interaction variability generically lose their stability through Hopf bifurcations. Limit cycles emerge, where surviving species have periodically oscillating abundances. At still stronger interactions, strange attractors appear, possibly from cascades of bifurcations, which lead to a chaotic dynamics of population abundances characterized by a positive largest Lyapunov exponent. This route to chaos illustrates how stationarity, oscillating periodicity and chaos in the dynamics of species abundances exist in rather general models of population dynamics, depending on $\sigma$ and $\gamma$. One important result is that all observed population dynamics in multi-species ecosystems can be reproduced by a unified and parcimonious mathematical model.
@@
Much of the existing literature on theoretical ecology is based on the surmise that the observed states and dynamics of ecosystems can be described by the time-asymptotic behavior of mathematical models. The latter is characterized by either a time-independent, stationary distribution of population abundances, or by persistently varying abundances, oscillating periodically or aperiodically. Accordingly, there has been a significant focus on the equilibrium fixed points of large systems of competing species, their stability and feasibility. May (1972); Roberts (1974); Rieger (1989); Allesina and Tang (2012); Tang, Pawar, and Allesina (2014); Fyodorov and Khoruzhenko (2016); Clenet, Massol, and Najim (2023); Ros et al. (2023) One key parameter is the variability $\sigma$ of the interspecies interaction. It has been found that a unique, asymptotically stable fixed point exists at weak enough $\sigma<\sigma_{c}$May (1972); Rieger (1989); Clenet, Massol, and Najim (2023); Ros et al. (2023) and that $\sigma_{c}$ increases in ecosystems with dominating predator-prey interactions. Allesina and Tang (2012); Tang, Pawar, and Allesina (2014) For $\sigma>\sigma_{c}$, one enters a phase with multiple unstable equilibria. Ros et al. (2023) Stabilization may still occur via species extinction, which drives the ecosystem to a novel, stable fixed point with reduced biodiversity. Beyond fixed points, the emergence of limit cycles through a Hopf bifurcation has been emphasized, albeit for small systems with only few species. Gilpin (1975); Coste, Peyraud, and Coullet (1979); Gardini, Lupini, and Messia (1989); Fussmann et al. (2000); Mcgehee et al. (2008) Further regimes with aperiodic persistent dynamics have recently been highlighted at stronger interactions and/or larger number of species, Hu et al. (2022) in ecosystems with sparse interactions, Roy et al. (2020); Marcus, Turner, and Bunin (2024) or communities with migrations. Dalmedigos and Bunin (2020)
In contrast to this mainstream philosophy focusing on the time-asymptotic behavior of mathematical models, several works have advocated that sufficiently large ecosystems have long relaxation times so that, by nature, they are always observed in a transient state. Hastings and Higgins (1994); Schreiber (2003); Morozov, Banerjee, and Petrovskii (2016); Hastings et al. (2018) Such large relaxation times may follow for instance from the slowing down of the dynamics in ecosystems close to criticality. Solé et al. (2002); Mora and Bialek (2011) Below we show that all these different behaviors naturally emerge in different regimes of a single mathematical model, as a function of only two parameters.
II Model and method
@@
Population dynamics in multi-species ecosystems is commonly studied in the framework of the generalized Lotka-Volterra model. Goel, Maitra, and Montroll (1971) Here we consider the variant investigated in Ref. Ros et al., 2023, which reads
@@
Equation (1) determines the time-evolution of the normalized abundance $N_{i}(t)\geq 0$ of species $i=1,2,\ldots S$, as a function of its intrinsic growth rate $\kappa_{i}$ and its interaction with other species. Interspecies interactions have a finite average $\mu/S$ and a variability $\sigma/\sqrt{S}$. The chosen scaling of interaction strengths with $S$ guarantees that the spectral support of the stability matrix [see Eq. (3b) below] does not change with $S$. Formulating the system as in Eq. (1) emphasizes the distribution underlying the random inter-species interactions. It allows us to pinpoint the relevant parameters in the route to chaos.
Fluctuations in interaction strengths between pairs of species are encoded in the components ${\mathbb{A}}_{ij}$ of the interaction matrix $\mathbb{A}$. Being interested in large, heterogeneous ecosystems with no particular structure, we follow a random matrix theory (RMT) approach, May (1972); Allesina and Tang (2015); Akjouj et al. (2024) where the matrix $\mathbb{A}$ belongs to an ensemble of matrices with elements ${\mathbb{A}}_{ij}$ that are normally distributed with vanishing ensemble average, $\langle{\mathbb{A}}_{ij}\rangle=0$, and covariances given by
@@
The cross-diagonal covariance parameter $\gamma\in[-1,1]$ allows to tune the types of two-species interactions that are likely to be encountered in the model. For $\gamma=1$, the interaction matrix $\mathbb{A}$} is totally symmetric and therefore interactions are either mutualistic (${\mathbb{A}}_{ij}\leq 0$ and ${\mathbb{A}}_{ji}\leq 0$), or competitive (${\mathbb{A}}_{ij}\geq 0$ and ${\mathbb{A}}_{ji}\geq 0$). For $\gamma=-1$, on the other hand, $\mathbb{A}$} is totally antisymmetric and all interactions are predator-prey (with ${\mathbb{A}}_{ij}\leq 0$ and ${\mathbb{A}}_{ji}\geq 0$ of vice-versa). Tuning $\gamma$ between $1$ and 1-1, one goes from a system dominated by mutualistic or competitive interaction ($\gamma$ close to $1$) to a system dominated by predator-prey interactions ($\gamma$ close to 1-1). gam Even though the mean $\mu$ introduces a bias towards mutualistic/competitive interactions, its contribution becomes negligible with respect to the contribution of the variability $\sigma$ for large $S$, due to the respective scalings in $S$ of the corresponding terms in Eq. (1). For $\gamma=0$, $\mathbb{A}$ belongs to the Ginibre ensemble or random matrices. Ginibre (1965)
The generalized Lotka-Volterra model of Eq. (1) assumes that $N_{i}(t)$ is real and varies continuously. This is a legitimate assumption only as long as $N_{i}(t)$ is sufficiently large. Large fluctuations have been observed in numerical simulations of Eq. (1), where some species resurrect from minuscule abundances, effectively corresponding to extinction. As a matter of fact, the dynamics of Eq. (1) leads to species extinction only for asymptotically long times, and the standard procedure to solve this atto-fox problem Mollison (1991) is to introduce a small, but finite extinction threshold $N_{c}$. When $N_{i}(t_{c})<N_{c}$, extinction occurs and $N_{i}(t>t_{c})\equiv 0$. We will set $N_{c}=10^{-20}$ but have checked that other choices lead to the same conclusions as presented below.
The model of Eq. (1) has been the focus of many recent investigations in theoretical ecology, yet, it is important to keep in mind the assumptions on which it is based. First, it neglects spatial structures present in real ecological networks. Second, it assumes that interspecies interactions are randomly distributed and neglects higher-order interactions involving three or more species. Third, and finally, a linear intrinsic growth rate $\kappa_{i}N_{i}$ is assumed, whereas some empirical data suggest a sublinear growth rate, which may have a significant impact on stability vs. biodiversity. Hatton et al. (2024); Aguadé-Gorgorió et al. (2025)
From now on, our focus is on the parameters $\sigma$ and $\gamma$ defining the variability of the interspecies interactions and accordingly we fix the growth rates, the initial number of species and the average interaction at values $\kappa_{i}\equiv 1$, $S=157$ and $\mu=5$. We have checked that varying $\mu$ does not change our conclusions as long as $\mu>0$.
@@
The RMT approach to ecosystem dynamics dates back at least to May’s seminal work. May (1972) Fixed-points $\vec{N^{*}}$ of Eq. (1) are defined by $\dot{N}_{i}^{*}=0$ and the dynamics in their vicinity is governed by the stability matrix $\mathbb{M}$
@@
$\mathbb{M}_{ij}$ $=$ $-N_{i}^{*}\left(\delta_{ij}+\frac{\mu}{S}+\frac{\sigma}{\sqrt{S}}{\mathbb{A}}_{ij}\right)\,.$ (3b)
[…]
$\sigma$ $<(1+\gamma)^{-1}\,.$ (4)
Beyond that border, fixed-point stability may be recovered via species extinctions, $N_{i}^{*}\rightarrow 0$, except for a number $N_{s}<S$ of surviving species, because then, the stability matrix $\mathbb{A}^{(r)}$ is reduced by removing rows and lines from $\mathbb{A}$, corresponding to the extinct species. Then, the spectrum of $\mathbb{A}^{(r)}$ has an elliptic support with semi-axes $\sigma(1\pm\gamma)\sqrt{N_{s}/S}$. Stability then bounds the number of surviving species as
$N_{s}/S\leq[\sigma(1+\gamma)]^{-2}\,.$ (5)
Note that this argument says nothing about the bifurcation through which fixed-points lose their stability.
@@ III Results
For $\gamma\neq 1$ the eigenvalues of the real, asymmetric matrix $\mathbb{M}$ are either real, or come in complex-conjugated pairs. This opens up the possibility that the fixed-point loses its stability through a Hopf bifurcation, after which the ecosystem dynamics is attracted to a stable limit cycle. Such a Hopf bifurcation is illustrated in Fig. 1, where the top panel shows periodic oscillations in abundances of the surviving species which are related in the bottom panel to the crossing of the imaginary axis by a complex-conjugated pair of eigenvalues. For large random matrices with $\gamma=0$ it has been shown that only ${\cal O}(\sqrt{S})$ of the $S$ eigenvalues are real, Edelman, Kostlan, and Shub (1994); Edelman (1997); Kanzieper and Akemann (2005) which suggests that such Hopf bifurcations are the rule rather than the exception in ecosystems with high biodiversity.
[…]
The prevalence of limit cycles depends on both the variability $\sigma$ of interspecies interaction and on the off-diagonal covariance parameter $\gamma$. As a matter of fact, to have a limit cycle, one needs a large enough $\sigma$ to have a bifurcation in the first place, moreover, the bifurcation must also involve a pair of complex-conjugated eigenvalues of the stability matrix. Both the value of $\sigma$ and the probability that a complex-conjugated pair of eigenvalues triggers the bifurcation depend on $\gamma$. One expects that more negative values of $\gamma$, together with the associated larger value of $\sigma$ to have a bifurcation favor the occurence of limit cycles. This is confirmed in Fig. 3, which shows the probability $P(\gamma)$ that the ecosystem has reached a state of persistent dynamics for given values of $\sigma$ and $\gamma$. Fig. 3 also confirms that for $\gamma<\sigma^{-1}-1$, below the threshold given by Eq. (4), one is in the regime with a unique asymptotically stable fixed point, without persistent dynamics. The numerical detection method employed for this statistical analysis cannot differentiate between limit cycles and chaotic motion on a strange attractor, because this would require to evaluate the largest Lyapunov exponent in each case, which is computationally time-consuming. However sampling those data indicate that chaotic motion sets in only at larger interactions, and constitutes a significant fraction of the data only for $\sigma=4$. Work to better quantify this fraction is currently underway and goes beyond the scope of this article. Data for larger $\sigma$ are not shown as they exhibit mass extinctions, with only very few surviving species at best, except for the smallest values of $\gamma$.
Figure 3: Probability to find the system in a state of persistent dynamics – either a limit cycle or a state of chaotic dynamics on a strange attractor – as a function of $\gamma$, and for $\sigma=1.5$ [Panel(a)], 2 [(b)], 3 [(c)] and 4 [(d)]. Data have been calculated over ensembles of 100 different interaction matrices, each with 5 different initial conditions. The grey area indicates the range of $\gamma$ where the system converges toward a unique stable fixed point, according to Eq. (4).
The observed increased probability to find limit cycles at negative values of $\gamma$ is related to the associated increased fraction of pairs of complex-conjugated eigenvalues in the spectrum of the stability matrix. Neglecting inhomogeneities in the fixed-point abundances $N_{i}^{*}$, this spectrum is real for $\gamma=1$ and purely imaginary for $\gamma=-1$, and it is expectable that the fraction of real eigenvalues decreases as $\gamma$ decreases. To the best of our knowledge, the only mathematically rigorous result at intermediate values of $\gamma$ is that only a fraction ${\cal O}(\sqrt{S})$ of the eigenvalues are real for the Ginibre ensemble, i.e., at $\gamma=0$Edelman, Kostlan, and Shub (1994); Edelman (1997); Kanzieper and Akemann (2005) While these observations suggest more frequent occurences of limit cycles as $\gamma$ decreases, what truly matters is whether the extreme eigenvalue with largest real part is real or complex, for a fixed realization of the interaction matrix in Eq. (1) with the covariance of Eq. (2). Calculating how likely that is, as a function of $\gamma$, would be a formidable task and we are unaware of any rigorous result in this direction. We therefore resort to numerical calculations (see supplementary material, Fig. S1). As expected, we find that the probability of a complex extreme eigenvalue increases with decreasing $\gamma$. To translate this result into a probability to have a limit cycle emerge through a Hopf bifurcation, we still need to take into account that there is no bifurcation at $\gamma=-1$, since there, increasing $\sigma$ only stretches the spectrum in the imaginary direction. We therefore expect that moderately negative values of $\gamma$ should favor the emergence of limit cycles. This qualitative argument is confirmed by the data shown in Fig. 3. A second numerical result shown in Fig. 5 (see supplementary material) is that the probability that the extreme eigenvalue is complex increases with the number $S$ of species.
@@
We translate these results into the language of theoretical ecology. Noting that $\gamma$ interpolates between ecosystems consisting purely of pairs of either mutualistic or competitive species for $\gamma=1$, to ecosystems with only predator-prey pairs of species for $\gamma=-1$ gam , persistent oscillating behavior in population abundances are expected to be prevailing in large, sufficiently interacting ecosystems with a majority of predator-prey pairs. We note that the occurence of Hopf bifurcations is of importance for theoretical ecology. It extends the stability of population coexistence beyond the loss of fixed-point stability, without the need for species extinction. The resulting increasing biodiversity is illustrated in Fig. 4. It is obvious that the average number of coexisting species at fixed $\sigma$ is higher for ecosystems equilibrating to limit cycles ($\langle N_{s}\rangle=58.4$ and $63$) than to fixed-points ($\langle N_{s}\rangle=51.8$ and $53.7$).
IV Conclusions